OpenStudy (agl202):

I think it's D. What is the equation of the following graph in vertex form? y = (x - 2)2 + 1 y = (x - 1)2 + 2 y = (x + 2)2 + 1 y = (x + 2)2-1

OpenStudy (agl202):

It's acually A.

OpenStudy (jango_in_dtown):

see the vertex is (2,1) and this kind of parabola has the equation y-1=(x-2)^2 implies y=(x-2)^2+1

OpenStudy (jango_in_dtown):

@Agl202

OpenStudy (agl202):

Thx... A little to late thought... :)